If P(A) = 0.82, P(D) = 0.83, and P(D∩A) = 0.78, then P(A|D') is Blank 1 and P(D|A) is Blank 2.


4 Decimal Places


1
Expert's answer
2021-05-07T11:50:51-0400

"P(A) = 0.82, P(D) = 0.83,P(D\u2229A) = 0.78"



 "P(A|D)'=1-P(A|D)=1-\dfrac{P(A\cap D)}{P(D)}=1-\dfrac{0.78}{0.83}=1-0.9397=0.06"


"P(D|A)=\dfrac{D\cap A)}{P(A)}=\dfrac{0.78}{0.82}=0.9512"